### v1

# Effective Free Energy Landscapes and Hawking-Page Transitions

Abstract
A recent interesting development on the dynamics of black hole phase transitions, the Hawking-Page transition, has been the so-called Gibbs free energy landscape approach. In this formalism, it is assumed that there exists a canonical ensemble of a series of black hole spacetimes with arbitrary horizon radius at a given ensemble temperature. An off-shell Gibbs free energy is defined for every spacetime state in the ensemble, with the horizon radius treated as the order parameter. The minima (maxima) of this function correspond to the various stable (unstable) black hole states. This off-shell Gibbs free energy is then treated as a classical effective drift potential of an associated Fokker-Planck equation used to study the stochastic dynamics of black hole phase transition under thermal fluctuations. Additive noise, which is independent of the black hole size, is assumed in obtaining the Fokker-Planck equation. In this work we extend the previous treatment by considering the effects of multiplicative noise, namely, noise that could scale with black hole size. This leads to an effective free energy function that can be used to study the modification of the Hawking-Page transition of a black hole system. It is realized that it is generally difficult to form black holes under a multiplicative noise, unless the effective and the original free energy become extremal at the same horizon radius. For this latter situation some theoretical noise profiles which are monotonically increasing/deceasing functions of the horizon radius are considered. It is found that stronger noise disfavors the formation of black hole.
Introduction.– In the early seventies of the last century, Bekenstein and Hawking revealed the thermal nature of black holes by associating a black hole with an entropy and a temperature B ; H ; BBH . Since then black hole thermodynamics has been an exciting and intriguing area of research in black hole physics that establishes a deep link between gravity, thermodynamics, and quantum physics (for reviews, see eg., W ; C ).
Among various aspects of black hole thermodynamics, the study of structures and transitions of thermodynamic phases of some black hole systems has received great interest in recent years [6-16]. Hawking-Page phase transition was found for the Schwarzschild-AdS (anti-de Sitter) black hole system, which is a first order transition between a thermal AdS space and the large AdS black hole at a certain critical temperature HP . The first order phase transition between small and large black holes was studied for the charged Reissner-Nordström (RN)-AdS black hole CEJM ; Wu , which later was shown to resemble that of the van der Waals liquid-gas system D1 ; D2 ; M1 . These interesting results soon inspired various studies of phase structures and transitions of other black hole systems (for a review, see eg., M2 ).
A recent interesting development on the dynamics of black hole phase transitions has been the so-called free energy landscape approach LW1 ; LW2 ; LW3 . In this formalism, it is assumed that there exists a canonical ensemble of a series of black hole spacetimes with arbitrary horizon radius at a given ensemble temperature. This ensemble consists also intermediate states which are not solutions of the Einstein equation as well as stable and unstable black hole solutions. An off-shell Gibbs free energy, obtained by replacing the Hawking temperature of the on-shell free energy by the ensemble temperature, is defined for every spacetime state in the ensemble. The horizon radius is treated as the order parameter. The minima (maxima) of this function correspond to the various stable (unstable) black hole states. This off-shell Gibbs free energy is then treated as an effective drift potential of an associated Fokker-Planck equation used to study the stochastic dynamics of black hole phase transition under thermal fluctuations. This approach has been applied to phase dynamics of the the Schwarzschild-AdS systems in LW1 , and to the RN-AdS systems in LW2 . Subsequently, other forms of free energy landscapes have been proposed, such as the Landau free energy X1 and the thermal potential X2 . The landscape approaches have since been extended to other black holes systems Y ; LW4 ; LW5 ; W , and Kramers escape rates of phase transitions have also been considered recently A1 ; X3 ; A2 .
So far the study of the stochastic dynamics of the black hole phase transitions is based mainly on uniform additive thermal noise which is independent of the horizon radius. However, in real situations it is not unreasonable to assume that the noise would depend on the size of black hole, considering the fact that thermal, quantum, and spacetime fluctuations near black holes of different sizes could be different. In this work we would like to explore the effect on the Hawking-Page transition if the noise could scale with space. Such noise is called multiplicative noise.
Fokker-Planck equation.– Consider a four-dimensional Schwarzschild-AdS black hole with mass $M$ and AdS curvature radius $L=\sqrt{-3/\Lambda}$, where $\Lambda$ is the cosmological constant. The metric (in $G=1$ units) is LW1
$ds^{2}$ $=$ $-f(r)dt^{2}+f(r)^{-1}\,dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$
$f(r)$ $=$ $1-\frac{2M}{r}+\frac{r^{2}}{L^{2}}.$ (1)
Setting $f(r)$ to zero gives the black hole horizon radius $r_{+}$, which in this case there is only one solution. In terms of $r_{+}$, the mass $M$, the Hawking temperature $T_{H}$, and the Bekenstein-Hawking entropy $S$ are
$M=\frac{r_{+}}{2}\left(1+\frac{r_{+}^{2}}{L^{2}}\right),~T_{H}=\frac{1}{4\pi r_{+}}\left(1+\frac{3r_{+}^{2}}{L^{2}}\right),~S=\pi r_{+}^{2}.$ (2)
The expression of $T_{H}$ implies that a Schwarzschild-AdS black hole has a minimal temperature (c.f. Fig. 1)
$\mathcal{T}_{m}=\frac{\sqrt{3}}{2\pi L}.$ (3)
[…]
$\mathcal{G}(r_{+},\mathcal{T})=\frac{r_{+}}{2}\left(1+\frac{r_{+}^{2}}{L^{2}}\right)-\pi{\mathcal{T}}r_{+}^{2},$ (4)
or, in terms of the two dimensionless variables $r\equiv r_{+}/L$ and $T\equiv{\mathcal{T}}L$,
$\mathcal{G}(r_{+},\mathcal{T})$ $=$ $L\,G_{0}(r,T),$ (5)
$G_{0}(r,T)$ $=$ $\frac{r}{2}\left(1+r^{2}\right)-\pi Tr^{2}.$
The minima (maxima) of this function correspond to the various stable (unstable) black hole states, as $G_{0}^{\prime}=0$ gives the relation between $T_{H}$ and $r_{+}$ in (2).
The Hawking-Page transition is easily understood in this landscape picture using $G_{0}(r,T)$, as shown in Fig. 2. There are two critical temperatures: the minimal temperature $T_{m}$ in (3), where $G_{0}$ exhibits an inflection point at $r=1/\sqrt{3}$, and the Hawking-Page temperature $T_{HP}=1/\pi$, where $G_{0}$ has two degenerate global minima at $r=0,1$. For $T<T_{m}$, there is just one global minimum of $G_{0}$ at $r=0$, representing the system is in a pure radiation phase, or the thermal AdS space. When $T_{m}<T<T_{HP}$, a local maximum and a local minimum appear for $r>0$, corresponding to an unstable small black hole phase and a metastable large black hole phase, respectively. For $T>T_{HP}$, the large black hole phase is the stable state.
The stochastic kinetics of the states in a thermodynamic ensemble under thermal fluctuation can be developed in terms of a Langevin equation with $\mathcal{G}$ as the external force potential, and a stochastic noise $\xi(t)$, which in LW1 is assumed implicitly to be a space-independent Gaussian noise,
$\dot{r}_{+}(t)=-\frac{\mathcal{G}^{\prime}(r_{+},\mathcal{T})}{\zeta}+\xi(t).$ (6)
Here $\zeta$ is the dissipation coefficient, the “dot” and the “prime” represent derivatives with respect to time $t$ and space $r_{+}$, respectively, and
$\langle\xi(t)\rangle=0,~~\langle\xi(t)\xi(t^{\prime})\rangle=2D\delta(t-t^{\prime}),~~~D\equiv\frac{k\mathcal{T}}{\zeta},$ (7)
where $k$ is the Boltzmann constant. The Fokker-Planck equation corresponding to (6) is R
$\frac{\partial}{\partial t}P(r_{+},t)=\frac{\partial}{\partial r_{+}}\left(\frac{\mathcal{G}^{\prime}(r)}{\zeta}\,P(r_{+},t)\right)+D\frac{\partial^{2}}{\partial r_{+}^{2}}\,P(r,t).$ (8)
Here $P(r,t)$ is the probability density function of the states in the thermal ensemble. The stationary state is
$P_{0}(r_{+})$ $\sim$ $\exp\left(-\frac{\mathcal{G}(r_{+},\mathcal{T})/\zeta}{D}\right)$
$\sim$ $\exp\left(-\frac{\mathcal{G}(r_{+},\mathcal{T})}{k\mathcal{T}}\right).$
Using (8), one can then study various aspects of the stochastic kinetics of the thermal ensemble of the black hole states, such as the mean first passage time for the black hole state switching and Hawking-Page transition LW1 .
Effective free energy landscape.– In (6) the noise is additive, as it is space-independent and added directly to the equation governing the change in $r$. But in real situations it is not unreasonable to assume that the noise would depend on the size of black hole, considering the fact that thermal, quantum, and spacetime fluctuations near black holes of different sizes could be different. As such, we would like to explore the effect on the Hawking-Page transition if the noise could scale with space. Such noise is called multiplicative noise.
The general Langevin equation with such a noise force is
$\dot{r_{+}}(t)=-\frac{\mathcal{G}^{\prime}(r_{+},T)}{\zeta}+g\left(\frac{r_{+}}{L}\right))\xi(t).$ (10)
Here the noise profile $g(\cdot)\neq 0$ is a dimensionless scaling function of the Gaussian noise, which for simplicity we assume to be time and temperature independent. The Fokker-Planck equation corresponding to (10) is R
$\frac{\partial}{\partial t}P=\frac{\partial}{\partial r_{+}}\left[\left(\frac{\mathcal{G}^{\prime}}{\zeta}-\lambda\,Dgg^{\prime}\right)P\right]+\frac{\partial^{2}}{\partial r_{+}^{2}}\left(Dg^{2}P\right).$ (11)
Here the parameter $\lambda=0,1$ according to Itô’s and Stratronovich’s rule, respectively, of converting a stochastic Langevin equation into a corresponding Fokker-Planck equation R .
The stationary state is
$P_{0}(r_{+})$ $\sim$ $\exp\left(-\frac{G(r_{+},\mathcal{T})}{k\mathcal{T}}\right)$ (12)
$G(r_{+},\cal{T})$ $\equiv$ $\int^{r_{+}}_{c}\,\frac{\mathcal{G}^{\prime}}{g^{2}}dr_{+}+k\mathcal{T}\left(2-\lambda\right)\ln g,$
where $c$ is a constant. Comparing with (Effective Free Energy Landscapes and Hawking-Page Transitions), this solution is the same as the stationary state of a Fokker-Planck equation (8) with $G$ as the drift potential. Thus the phase structure of the black hole states in the thermodynamic ensemble with multiplicative noise can be equivalently studied as that with an additive noise, but with $G$ as the effective Gibbs free energy landscape.
In terms of the dimensionless variables, $G(r_{+},\mathcal{T})$ is
$G(r,T)$ $=$ $\int^{r}_{c}\,\frac{1}{g^{2}}\left(\frac{3}{2}r^{2}-2\pi Tr+\frac{1}{2}\right)dr$ (13)
$+\frac{kT}{L^{2}}\left(2-\lambda\right)\ln g(r).$
For a uniform additive noise $g(r)=1$, we have $G(r,T)=G_{0}(r,T)$. So the classical landscape remains unchanged. This is not the case for general noise profile. Particularly, unlike the case with $G_{0}(r,T)$, the extrema of $G(r,T)$ do not necessarily correspond to the black holes solutions of the Einstein equations for general noise, unless $G^{\prime}(r,T)=0$ and $G_{0}^{\prime}(r,T)=0$ share same common roots at a given temperature $T$ (at these common roots we have $g^{\prime}(r)=0$). Thus it is seen that it is generally difficult to form black holes under a multiplicative noise.
If $G(r,T)$ has the same functional form as $G_{0}(r,T)$ in some part of the half-line in which $G_{0}(r,T)$ has extrema, then $G(r,T)$ admits black hole solutions. For instance, one might consider the situation where thermal fluctuations differ significantly from additive white noise only for black hole states with small horizon radii. This is not unreasonable in view of the strong gravitational and quantum effects near small black holes.
Here we would like to consider one such theoretical noise profile to see how multiplicative noise could affect the nature of the Hawking-Page transition of a Schwarzschild-AdS black hole systems. For definiteness, we adopt the Stratronovich’s rule ($\lambda=1$) and set $k=L=1$. The Itô’s rule ($\lambda=0$) gives qualitatively similar results.
We take $g(r)=1+a-sign(a)\sqrt{a^{2}-(r-|a|)^{2}}$ for $r<|a|$, and $g(r)=1$ for $r>|a|$ (a>1a>-1). The noise differs from the uniform additive noise only in $r<|a|$, and it is stronger (weaker) than the additive white noise for $a>0$ (1<a<0-1<a<0). For $a>0$, $g(r)$ decreases from $g(0)=1+a$ to $g(a)=1$ along a circular arc of radius $a$. For $a<0$, $g(r)$ increases from g(0)=1|a|g(0)=1-|a| to $g(a)=1$ along a circular arc. With this profile, black hole spacetimes with smaller horizons experience greater changes of noise level from the the additive noise in (5).
Fig. 3 shows $G(r,T_{HP})$ for different values of $a$ at $T_{HP}=1/\pi$, the Hawking-Page critical temperature in the case of additive noise (i.e., $a=0$). The constant $c$ is chosen such that $G(0,T)=0$. It is seen that for 1<a<0-1<a<0 (weaker noise level near $r=0$), the system is still in the thermal AdS state, and for $a>0$ (stronger noise level near $r=0$), the system is already in the large black hole state. One can understand this effect as weaker noise fluctuation near $r=0$ favors the formation of the thermal AdS states and disfavors the formation of the large black hole. Thus one would expect a higher (lower) Hawking-Page critical temperature for 1<a<0-1<a<0 ($a>0$). It is worthy to note that for $a>0$, $G(r,T)$ could develop a new local minimum in $r<a$. However this minimum does not correspond to a solution of the Einstein equation.
RN-AdS black hole.– We now extend the previous consideration to the corresponding situation for the RN-AdS system as discussed in LW2 .
Following LW2 , the basic data of RN-AdS system are summarized below. The metric in this case is given by (1) but with $f(r)$ changed to
$f(r)=1-\frac{2M}{r}+\frac{r^{2}}{L^{2}}+\frac{Q^{2}}{r^{2}},$ (14)
where $Q$ is the charge of the black hole. The Schwarzschild-AdS black hole can be viewed as the $Q\to 0$ limit of the RN-AdS system.
The largest root of the equation $f(r)=0$ gives the radius $r_{+}$ of the event horizon. The mass $M$ and the Hawking temperature $T_{H}$ are
$M$ $=$ $\frac{r_{+}}{2}\left(1+\frac{r_{+}^{2}}{L^{2}}+\frac{Q^{2}}{r_{+}^{2}}\right),$ (15)
$\mathcal{T}_{H}$ $=$ $\frac{1}{4\pi r_{+}}\left(1+\frac{3r_{+}^{2}}{L^{2}}-\frac{Q^{2}}{r_{+}^{2}}\right).$ (16)
In LW2 thermal phase structure of the RN-AdS black hole system at different ensemble temperature $\mathcal{T}$ is studied by treating the cosmological constant $L$ (in terms of the related thermal pressure $P=3/8\pi L^{2}$ D1 ) as a control parameter, keeping the charge $Q$ fixed. However, in order to be in line with the previous discussion, we will keep $L$ fixed (and set $L=1$ for numerical analysis) and treat $Q$ as the control parameter. Thus we will rewrite some relevant formulas in LW2 for our purpose.
In terms of the dimensionless variables $r=r_{+}/L,T=\mathcal{T}L$ and $q=Q/L$, The Hawking temperature can be re-expressed as
$T_{H}=\frac{1}{4\pi r}\left(1+3r^{2}-\frac{q^{2}}{r^{2}}\right).$ (17)
$T_{H}$ is a monotonic function of $r$ when $q>q_{c}=1/6$, and has a local minimum $T_{m}$ and a local maximum $T_{M}$ otherwise. Fig. 4 depicts these situations for $q=1/3$ and $1/10$.
It is clear that there is no phase transition if $q>q_{c}$. For $q<q_{c}$, $T_{m}$ and $T_{M}$ are given by
$T_{m/M}=\frac{1}{\pi}\sqrt{\frac{3}{2}}\frac{\left(1-12q^{2}\pm\sqrt{1-36q^{2}}\right)}{\left(1\pm\sqrt{1-36q^{2}}\right)^{3/2}}.$ (18)
Note that in the limit $q\to 0$, $T_{M}$ does not exist and $T_{m}$ reduces to (3) (in dimensionless form). For $T_{m}<T_{H}<T_{M}$ there exists three branches of black hole solutions: small, large and (unstable) intermediate black holes. There can be a first-order phase transition between the small and large black holes.
As before, one considers a canonical ensemble at a temperature $T$ consisting of a series of black hole spacetimes with arbitrary horizon radius. Spacetimes other than the three branches of the black holes are off-shell, i.e., they are not solutions of the Einstein equation. The off-shell Gibbs free energy function is
$G_{0}(r,T)=\frac{r}{2}\left(1+r^{2}+\frac{q^{2}}{r^{2}}\right)-\pi Tr^{2}.$ (19)
In Fig. 5 we give the graphs of $G_{0}(r,T)$ with $q=1/10$ for different range of $T$. Here $T_{m}=0.27120$ and $T_{M}=0.34869$. For $T<T_{m}$, $G_{0}$ has one global minimum corresponding to the small black hole. An inflection point occurs at $T_{m}$. For $T$ in between $T_{m}$ and $T_{M}$ there are two local minima and a local maximum corresponding to the three black holes solutions mentioned before. When $T>T_{M}$, there is only one global minimum for the large black hole. Hawking-Page transition between the small and large black holes occurs at $T_{HP}=0.28475$.
As mentioned before, the effective Gibbs free energy is still given by (19) in the presence of an additive noise. For multiplicative noise, one finds from (13)
$G(r,T)$ $=$ $\int^{r}_{c}\,\frac{1}{g^{2}}\left(\frac{3}{2}r^{2}-2\pi Tr+\frac{1}{2}-\frac{q^{2}}{2r^{2}}\right)dr$ (20)
$+\frac{k}{L^{2}}T\left(2-\lambda\right)\ln g(r).$
Phase structures and transitions can then be studied with this effective Gibbs function for different noise profile $g(r)$ as done in previous sections.
Again we consider only $g(r)=1+a-sign(a)\sqrt{a^{2}-(r-|a|)^{2}}$ for $r<|a|$, and $g(r)=1$ for $r>|a|$ (a>1a>-1). Fig. 6 shows $G(r,T_{HP})$ for different values of $a$ at $T_{HP}=0.28475$ (with $L=1,c=0.005$). The constant $c$ is chosen such that $G(0.05,T)=0$. We see that the system is still in the small black hole phase when $a<0$, and has already transited to the large black hole phase when $a>0$. This conforms to the previous observation that strong noise level disfavors black hole formation.
Summary.– In this work we have attempted to extend the previous work on Gibbs free energy landscape by considering the effects of multiplicative noise. An effective free energy function is obtained that can be used to study the modification of the Hawking-Page transition of a black hole system. It is realized that it is generally difficult to form black holes under a multiplicative noise, unless the effective and the original free energy become extremal at the same horizon radius. The latter situation was modeled by some theoretical noise profiles which are monotonically increasing/deceasing functions of the horizon radius. It is found that stronger noise disfavors the formation of black hole.
The work is supported in part by the Ministry of Science and Technology (MOST) of the Republic of China under Grants NSTC 113-2112-M-032-010 and NSTC 114-2112-M-032-005.